OpenStudy (anonymous):

Let C be parametrized by x = sin^2 t and y = sin t cos t for 0 ≤ t ≤ π/2. Find the length S of the curve C.

OpenStudy (anonymous):

\[dS = \sqrt{x'(t)^2 + y'(t) ^2 }dt\] integrate both sides ^_^

OpenStudy (anonymous):

... with limits \(\displaystyle\int_0^{\pi/2}\cdots\)

OpenStudy (anonymous):

@SithsAndGiggles got an imaginary answer 0.602343+0.381073 i and its wrong. x'=sin(2t) and y'=cos(2t) right?

OpenStudy (anonymous):

yes. remeber the ID:\[\sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

yeah but then I still dont know what I got wrong?